OpenStudy (barrelracing):

Please help so confused :( 77 is 55% if what?

OpenStudy (anonymous):

to get the answer you could set up a proportion \[\frac{ 77 }{ x }=\frac{ 55 }{ 100 }\]

OpenStudy (anonymous):

And solve for x

OpenStudy (anonymous):

Minus 5%. from 77. n x 2 = answer

OpenStudy (anonymous):

n = 77 - 5%

OpenStudy (barrelracing):

so do i cross multiply?

OpenStudy (anonymous):

Yes @Barrelracing

OpenStudy (anonymous):

That won't work @Syntex because 5% is proportionate not an exact value

OpenStudy (anonymous):

Is this a multiawnser question?

OpenStudy (barrelracing):

i got 140

OpenStudy (anonymous):

Yup that's right :)

OpenStudy (anonymous):

Was about to say that it would be around 140. kthx. bye. have a good day.

OpenStudy (barrelracing):

thank you everyone

OpenStudy (qwertty123):

Hey @cmtboy2 Sorry I coouldn't help you @Barrelracing